An arithmetic sequence has initial term 6 and common difference 624. A geometric sequence has initial term 2 and common ratio 3. Determine an n so the nth term of the arithmetic sequence is that same as the nth term of the geometric sequence.
I'm not sure how this works, and the only hint I was given was that initial term = a0. Any help would be so great, thanks for your time. (:

Answer by scott8148(6628)

6 + 624(n - 1) = 2[3^(n - 1)]

624n - 618 = (2/3) 3^n

936n - 927 = 3^n

you can find n by substituting values


HINT: between 6 and 10